# Sampling Theory for Function Approximation with Numerical Redundancy

### v1

@@ 5 Applications
5.1 Approximating with known asymptotic behaviour
An approximation problem typically stems from an underlying problem in science or engineering, where expert knowledge on its solution is often available. It is, however, non-trivial to incorporate this knowledge into an efficient numerical scheme. One way to do this is by adding extra basis functions to the spanning set, which capture the known properties of the function to be approximated. This forces us to compute with non-standard bases that are often numerically redundant. For instance, suppose our goal is to approximate
@@ 5.1 Approximating with known asymptotic behaviour
assuming $n$ is even. The functions $\varphi_{i}$ are standard (smooth) basis functions, while the singular behaviour is represented by the weight function $w$. For $w\in L^{\infty}(-1,1)$, this spanning set is a subsequence of an overcomplete linearly independent frame with frame bounds $A=1+\operatorname*{ess\,inf}_{x\in(-1,1)}\lvert w(x)\rvert^{2}$ and $B=1+\operatorname*{ess\,sup}_{x\in(-1,1)}\lvert w(x)\rvert^{2}$ [5, Example 3]. Hence, following section 2.1, $\Phi_{n}$ is numerically redundant for sufficiently large $n$. For further analysis and usecases of this type of spanning sets we refer to [5] and references therein. For the approximation of $f$ (56), we choose $\varphi_{i}$ equal to the Legendre polynomial of degree $i-1$ and $w(x)=\sqrt{(x+1)/2}$. In this case, $\Phi_{n}$ is a subsequence of a frame with frame bounds $A=1$ and $B=2$.
5.1.1 Properties of $\boldsymbol{\Phi_{n}}$
@@
On Figure 3, the approximation of $f$ (56) for three different spanning sets is shown: Legendre polynomials, monomials, and the Legendre + Weighted Legendre spanning set defined by (57). Figures 3(a) and 3(b) display the first eight elements of the spanning sets along with the singular value profile of their associated synthesis operator for $n=100$. The Legendre basis functions are orthonormal, such that all singular values of the synthesis operator are equal to $1$. In contrast, the monomials and the Legendre + Weighted Legendre spanning set are both numerically redundant for sufficiently large $n$. However, a clear qualitative difference exists between these two sets. As can be seen on Fig. 3(a), the monomials look increasingly more alike as their degree increases, such that high-degree monomials are indistinguishable when using finite precision. As a result, all singular values of the synthesis operator are exponentially decaying. The Legendre + Weighted Legendre spanning set contains half of the Legendre basis augmented with weighted Legendre functions. Hence, in contrast to the monomial basis, a large portion of the set is well-distinguishable. As a result, the synthesis operator has many singular values $\sigma_{i}\approx 1$, combined with exponentially decaying singular values. This matches the singular value profile described in section 4.3.
@@
On Fig. 3(c) and 3(d), the $L^{2}$-norm of the approximation error and the $L^{2}$-norm of the expansion coefficients are displayed for the approximation of $f$ using a truncated singular value decomposition (TSVD) solver with $\epsilon=10\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}$, where $\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}$ denotes the working precision associated with IEEE double precison floating point numbers. The approximations are computed using $5000$ Legendre points. Both the Legendre basis and the monomials span the space of polynomials up to degree $n-1$ and, therefore, theoretically exhibit the same algebraic rate of convergence for the approximation of $f$. However, the expansion coefficients of the monomial approximations blow up, such that the error of the numerically computed approximation stagnates, following the results of section 2.2. On the other hand, the approximation in the Legendre + weighted Legendre spanning set converges exponentially. The norm of the expansion coefficients initially grows, yet expansions with modest coefficient exist as $n\to\infty$, guaranteeing accurate numerical approximations. This is to be expected, since $\Phi_{n}$ is a subsequence of an infinite-dimensional linearly independent frame, see e.g. Theorem 4.
@@
5.1.2 Deterministic sampling
@@
The two-parameter plots on Figure 4 show $1/\sqrt{A_{m,n}}$, $1/\sqrt{A^{\epsilon}_{m,n}}$, and the uniform approximation error of the TSVD approximation to $f=J_{1/2}(x+1)+1/(x^{2}+1)$ in $\Phi_{n}$ (57) with $n=80$, for a varying number of Legendre points and exponentially clustered sample points. The regularization parameter equals $\epsilon=10\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}$, where $\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}$ denotes the working precision associated with IEEE double precison floating point numbers. From Figure 3(c), we know that the best approximation to $f$ in the Legendre + weighted Legendre spanning set is of the order of machine precision. Following Theorem 1 and Theorem 2, both $1/\sqrt{A_{m,n}}$ and $1/\sqrt{A^{\epsilon}_{m,n}}$ give an indication of the error of the discrete least squares approximation. From the numerical results, we conclude that $A_{n,m}^{\epsilon}$ accurately predicts the behaviour of the uniform approximation error, while $A_{n,m}$ overestimates the amount of samples that are needed for accurate numerical approximation.
Figure 4: Analysis of the Legendre + weighted Legendre spanning set defined by (57) for $n=80$. Left: $1/\sqrt{A_{n,m}}$, right: $1/\sqrt{A_{n,m}^{\epsilon}}$ for $\epsilon=10\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}\approx 2\times 10^{-15}$, bottom: the uniform approximation error of the TSVD approximation of $f(x)=J_{1/2}(x+1)+1/(x^{2}+1)$ with truncation threshold $\epsilon$. The black squares correspond to $A_{n,m}=0$. The white lines mark when the total number of samples equals $n=80$ or $n^{\epsilon}\approx 54.8$, the effective degrees of freedom. We conclude that $A_{n,m}^{\epsilon}$ accurately predicts the behaviour of the approximation error, while $A_{n,m}$ overestimates the amount of samples that are needed for accurate numerical approximation.
@@ 5.2 Approximating on irregular domains
Approximating a function on an irregular domain $D\subset\mathbb{R}^{d}$ is required for many computational problems, yet an orthonormal or Riesz basis is generally unknown for such a domain. Therefore, numerical methods often resort to working with an orthonormal basis defined on a bounding box surrounding $D$. Crucially, this set of functions does not form an orthonormal basis for the underlying domain $D$, but it does constitute an overcomplete frame, referred to as an extension frame. As discussed in section 2.1, a subsequence of an overcomplete frame inevitably becomes numerically redundant as its size increases and, hence, regularization is needed for numerical stability. Approximating on an irregular domain is, therefore, a typical scenario in which numerical redundancy arises. An alternative strategy is to compute with a numerically orthogonalized spanning set. However, section 6 shows that straightforward techniques to orthogonalize do not succeed in circumventing the effects of finite precision, and result in approximations that behave similarly as regularized approximations.
@@
for odd $n$. Smooth functions, not necessarily periodic on $[-W,W]$, can be approximated in $V_{n}=\operatorname{span}(\Phi_{n})$ with exponential convergence [35]. As mentioned before, regularization is needed for numerical stability, and lowers the accuracy of the approximation. The numerical convergence behaviour is extensively analysed in [7].
In this section, we analyse how regularization affects the required amount of data for accurate approximation using the results from section 4, which focus on random pointwise sampling. The associated analysis outlined in section 4.3 depends heavily on a good understanding of the singular values and the singular vectors associated to the synthesis operator $\mathcal{T}_{n}$. For many spanning sets, little is known about this; however, the singular vectors of the Fourier extension frame are related to discrete prolate spheroidal wave functions (DPSWFs), which have been analysed extensively in the field of signal processing. For a detailed explanation of this connection, see [38, section 3.1]. DPSWFs are an indispensable tool for bandlimited extrapolation, introduced and thoroughly analysed by Slepian and his collaborators during their time at Bell Labs [46]. In this setting, a notion of effective degrees of freedom is not new, see e.g. [37] and [21, section 2.3]
5.2.1 Effective degrees of freedom
In the case of pointwise random sampling, the required number of samples for accurate regularized least squares approximation scales with $n^{\epsilon}\log(n^{\epsilon})$, as stated in Theorem 10. Here, $n^{\epsilon}$ denotes the effective degrees of freedom. This theorem assumes that the samples are drawn randomly from a (near-)optimal sampling distribution defined by (46). The effective degrees of freedom $n^{\epsilon}$ thus play a similar role for regularized approximation as the dimension $n$ for unregularized approximation. The following theorem formalizes the dependence of $n^{\epsilon}$ on $n$, $W$ and $\epsilon$ for the one-dimensional Fourier extension frame.
@@ 5.2.1 Effective degrees of freedom
such that the effective degrees of freedom $n^{\epsilon}$ can be bounded by
@@
Note that $\sigma_{i}(\mathcal{T}_{n})^{2}$ equals the $i$-th eigenvalue of the Gram matrix $G_{n}$, representing the operator $\mathcal{T}_{n}^{*}\mathcal{T}_{n}$. Furthermore, $G_{n}$ is exactly the prolate matrix [36, eq. (1)]. Therefore, $\sigma_{i}(\mathcal{T}_{n})^{2}=\lambda_{i-1}$, for all $i=1,\dots,n$, where the latter is analysed in [36]. It follows from [36, Corollary 1] that (52) holds for (61). Moreover, (62) can be deduced by [36, Corollary 2]. The final result is a consequence of Theorem 11. ∎
A numerical verification of Theorem 12 is shown on Figure 5 for $n=141$ and $\epsilon=10\kern 1.0pt\operatorname{\epsilon_{\text{mach}}^{\text{sp}}}\approx 10^{-6}$ and $\epsilon=10\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}\approx 2\times 10^{-15}$, where $\operatorname{\epsilon_{\text{mach}}^{\text{sp}}}$ and $\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}$ denote the working precision associated with IEEE single and double precison floating point numbers, respectively. Note that the computation of $n^{\epsilon}$ is done in extended precision, such that the singular values $\sigma_{i}$ can be computed with a working precision significantly smaller than $\epsilon$. For implementational details, we refer to [31]. Both the theory and the numerical verification show that the effective degrees of freedom essentially scale as $2nW$. This formalizes the intuition that increased redundancy, i.e., smaller $W$, results in a reduced amount of required information. Observe that $n^{\epsilon}=\mathcal{O}(n)$, such that asymptotically the amount of samples needed for regularized and unregularized approximation is similar, namely $\mathcal{O}(n\log(n))$.
@@
Figure 5: The effective degrees of freedom $n^{\epsilon}$ for the one-dimensional Fourier extension frame (60) as a function of the domain parameter $W$, for $n=141$. The dashed lines illustrate $n^{\epsilon}\leq n=141$ and the theoretical bound $n^{\epsilon}\leq\lceil 2nW\rceil+2+\frac{2}{\pi^{2}}\left(\log(8/\epsilon^{2})+1\right)\log(4n)$ stated in Theorem 12.
@@ 5.2.2 Uniformly random sampling
For many applications, one cannot freely choose the sample distribution. Another interesting question to answer in this regard is how many uniformly random samples are needed for accurate approximation. As follows from Theorems 9 and 10 for $w=1$, this number is proportional to the maximum of the inverse Christoffel function $k_{n}$ and the inverse effective Christoffel function $k_{n}^{\epsilon}$ for unregularized and regularized least squares fitting, respectively. The following theorem shows that $\|k_{n}\|_{L^{\infty}[-W,W]}=\mathcal{O}(n^{2})$.
@@
For the spanning set defined by (60), the maximum of the inverse Christoffel function $k_{n}$ can be bounded by
$C_{1}\left(\left(1+\frac{\pi}{2}\right)n^{2}-\pi/2\right)\leq\|k_{n}\|_{L^{\infty}[-W,W]}\leq C_{2}\left(\left(1+\frac{\pi}{2}\right)n^{2}-\pi/2\right),$
where $C_{1},C_{2}>0$ are independent of $n$.
@@
An alternative characterization of the inverse Christoffel function [42, eq. (4.1.1)] is
$k_{n}(x)=\max_{v\in V_{n}}\frac{\lvert v(x)\rvert^{2}}{\|v\|^{2}_{L^{2}[-W,W]}},\qquad\text{such that }\qquad\|k_{n}\|_{L^{\infty}[-W,W]}=\max_{v\in V_{n}}\frac{\|v\|_{L^{\infty}[-W,W]}^{2}}{\|v\|_{L^{2}[-W,W]}^{2}}.$
From [26, Theorem 1] it immediately follows that there exist constants $C_{1},C_{2}>0$ such that
$C_{1}\left(\left(1+\frac{\pi}{2}\right)n^{2}-\pi/2\right)\leq\max_{v\in V_{n}}\frac{\|v\|^{2}_{L^{\infty}[-W,W]}}{\|v\|^{2}_{L^{2}[-W,W]}}\leq C_{2}\left(\left(1+\frac{\pi}{2}\right)n^{2}-\pi/2\right).$
The effective inverse Christoffel function $k_{n}^{\epsilon}$ can be bounded using (55), which depends on the size of the plunge region $i_{2}-i_{1}$ and $\operatorname{S_{\text{tail}}}$. The latter is shown to be $\mathcal{O}(\log(n))$ in Theorem 12. The following theorem proves that the size of the plunge region is $\mathcal{O}(\log(n))$ as well.
@@
$\|k_{n}^{\epsilon}\|_{L^{\infty}[-W,W]}\leq\frac{n}{1-\epsilon^{2}}+\left(4+\frac{2}{\pi^{2}}\left(2\log\left(\frac{8}{\epsilon^{2}}\right)+1\right)\log(4n)\right)\max_{i_{1}<i\leq n}\|u_{i}\|_{L^{\infty}[-W,W]}^{2}.$ (65)
@@
Similarly as for the proof of Theorem 12, observe that $\sigma_{i}(\mathcal{T}_{n})^{2}=\lambda_{i-1}$, for all $i=1,\dots,n$, where the latter is analysed in [36]. It follows from [36, Corollary 1] that (51) holds for (64). By combining the results from Theorem 12 with (55) and (64), we arrive at (65). ∎
It remains to examine how the basis functions $u_{i}$ behave. As mentioned before, these functions are linked to the discrete prolate spheroidal wave functions. In Appendix C, we summarize known asymptotic results due to Slepian, which indicate that $\|u_{i}\|_{L^{\infty}[-W,W]}=\mathcal{O}(\sqrt{n})$ as $n\to\infty$ for all $i$, such that
$\|k_{n}^{\epsilon}\|_{L^{\infty}[-W,W]}=\mathcal{O}(n\log(n)).$
This behaviour is verified numerically in Figure 6, which confirms that $\|k_{n}\|_{L^{\infty}[-W,W]}=\mathcal{O}(n^{2})$, while $\|k_{n}^{\epsilon}\|_{L^{\infty}[-W,W]}$ appears to grow linearly in $n$. This difference translates in the need for log-linear or quadratic oversampling when computing a least squares approximation with or without regularization, respectively, using uniformly random samples.
@@
The approximation seems to converge root-exponentially. For a more detailed analysis of the convergence behaviour, we refer to [7]. Note that convergence is exponential without regularization, see e.g. [35] and the example in section 6, but demands quadratic oversampling. Hence, we get root-exponential convergence as a function of the number of sample points $m$ with and without regularization. This is an example of a setting where the negative effects of regularization, i.e., reduced accuracy, are balanced by the positive effects, i.e., reduced need for data. What happens if we want to compute an approximation that converges root-exponentially all the way down to zero? In this case, we would need to vary the precision and the regularization parameter with $n$: $\epsilon\sim\epsilon_{\text{mach}}=\exp(-C\sqrt{n})$ for some $C>0$. Following Theorem 14, linear oversampling does not suffice in this case. Specifically, the required linear oversampling factor scales with $\log(1/\epsilon)$.
@@
Figure 6: The maximum of the inverse (effective) Christoffel function for the one-dimensional Fourier extension frame (60) as a function of $n$. Full line: $\|k_{n}\|_{L^{\infty}[-W,W]}$, circles: $\|k_{n}^{\epsilon}\|_{L^{\infty}[-W,W]}$ for single precision $\epsilon=10\kern 1.0pt\operatorname{\epsilon_{\text{mach}}^{\text{sp}}}\approx 10^{-6}$, and squares: $\|k_{n}^{\epsilon}\|_{L^{\infty}[-W,W]}$ for double precision $\epsilon=10\kern 1.0pt\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}\approx 2\times 10^{-15}$. The linear behaviour of $k_{n}^{\epsilon}$ versus the quadratic behaviour of $k_{n}$ implies a difference between log-linear and quadratic sampling rates for least squares fitting with uniformly random samples. The dashed lines mark $5\log_{10}(1/\epsilon)+C$, for varying constant $C$. They empirically match the behaviour of $k_{n}^{\epsilon}$.

### v2

@@ 6 Applications
6.1 Approximating with known singularities
An approximation problem typically stems from an underlying problem in science or engineering where expert knowledge on its solution is available. It is, however, non-trivial to incorporate this knowledge into an efficient numerical scheme. One way to do this is by adding extra basis functions to the spanning set, which capture the known properties of the function to be approximated. This forces us to compute with non-standard bases that are typically numerically redundant. We choose a simple example to illuminate the principles, noting that they hold much more generally.
Suppose our goal is to approximate
@@ 6.1 Approximating with known singularities
assuming $n$ is even. The functions $\varphi_{i}$ are standard (smooth) basis functions, while the singular behaviour is represented by the weight function $w$. For $w\in L^{\infty}(-1,1)$, this spanning set is a subsequence of an overcomplete linearly independent frame with frame bounds $A=1+\operatorname*{ess\,inf}_{x\in(-1,1)}\lvert w(x)\rvert^{2}$ and $B=1+\operatorname*{ess\,sup}_{x\in(-1,1)}\lvert w(x)\rvert^{2}$ [6, Example 3]. Hence, following section 2.1, $\Phi_{n}$ is numerically redundant for sufficiently large $n$. For further analysis and usecases of this type of spanning sets we refer to [6] and references therein. For the approximation of $f$ (60), we choose $\varphi_{i}$ equal to the Legendre polynomial of degree $i-1$ and $w(x)=\sqrt{(x+1)/2}$. In this case, $\Phi_{n}$ is a subsequence of a frame with frame bounds $A=1$ and $B=2$.
6.1.1 Numerical redundancy in action
@@
On Figure 6, the approximation of $f$ (60) for three different spanning sets is shown: Legendre polynomials, monomials, and the Legendre + Weighted Legendre spanning set defined by (61). Figures 6(a) and 6(b) display the first eight elements of the spanning sets along with the singular value profile of their associated synthesis operator for $n=100$. The Legendre basis functions are orthonormal, such that all singular values of the synthesis operator are equal to $1$. In contrast, the monomials and the Legendre + Weighted Legendre spanning set are both numerically redundant for sufficiently large $n$. However, a clear qualitative difference exists between these two sets. As can be seen on Figure 6(a), the monomials look increasingly more alike as their degree increases, such that high-degree monomials are indistinguishable when using finite precision. As a result, all singular values of the synthesis operator are exponentially decaying. The Legendre + Weighted Legendre spanning set contains half of the Legendre basis augmented with weighted Legendre functions. Hence, in contrast to the monomial basis, a large portion of the set is well-distinguishable. As a result, the synthesis operator has many singular values $\sigma_{i}\approx 1$, combined with exponentially decaying singular values. This matches the singular value profile described in section 5.3.
@@
On Figure 6(c) and 6(d), the $L^{2}$-norm of the approximation error and the $L^{2}$-norm of the expansion coefficients are displayed for the approximation of $f$ using a truncated singular value decomposition (TSVD) solver with $\epsilon=10\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}$, where $\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}$ denotes the working precision associated with IEEE double precison floating point numbers. The approximations are computed using $5000$ Legendre points. Both the Legendre basis and the monomials span the space of polynomials up to degree $n-1$ and, therefore, theoretically exhibit the same algebraic rate of convergence for the approximation of $f$. However, the expansion coefficients of the monomial approximations blow up, such that the error of the numerically computed approximation stagnates, following the results of section 2.2. On the other hand, the approximation in the Legendre + weighted Legendre spanning set converges exponentially. The norm of the expansion coefficients initially grows, yet expansions with modest coefficient exist as $n\to\infty$, guaranteeing accurate numerical approximations. This is to be expected, since $\Phi_{n}$ is a subsequence of an infinite-dimensional linearly independent frame, see e.g. Theorem 4.
This comparison demonstrates that numerical redundancy should not be viewed as a drawback; rather, allowing it enables the development of non-standard computational methods. The inclusion of weighted polynomials, for instance, leads to a significant improvement in the speed of convergence. However, not all numerically redundant sets are advantageous for computational purposes. For example, in the monomial basis no accurate solutions with bounded coefficients exist, resulting in a stagnation of the error of the best approximation in finite-precision arithmetic.
@@
6.1.2 Analysis of deterministic point sets
@@
The two-parameter plots on Figure 7 show $1/\sqrt{A_{m,n}}$, $1/\sqrt{A^{\epsilon}_{m,n}}$, and the uniform approximation error of the TSVD approximation to $f=J_{1/2}(x+1)+1/(x^{2}+1)$ in $\Phi_{n}$ (61) with $n=80$, for a varying number of Legendre points and exponentially clustered sample points. The regularization parameter equals $\epsilon=10\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}$, where $\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}$ denotes the working precision associated with IEEE double precison floating point numbers. From Figure 6(c), we know that the best approximation to $f$ in the Legendre + weighted Legendre spanning set is of the order of machine precision. Following Theorem 1 and Theorem 2, both $1/\sqrt{A_{m,n}}$ and $1/\sqrt{A^{\epsilon}_{m,n}}$ give an indication of the error of the discrete least squares approximation. We can conclude that $A_{n,m}^{\epsilon}$ accurately predicts the behaviour of the approximation error, while $A_{n,m}$ overestimates the amount of samples that are needed for accurate numerical approximation.
Figure 7: Analysis of the Legendre + weighted Legendre spanning set defined by (61) for $n=80$. Left: $1/\sqrt{A_{n,m}}$, right: $1/\sqrt{A_{n,m}^{\epsilon}}$ for $\epsilon=10\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}\approx 2\times 10^{-15}$, bottom: the uniform approximation error of the TSVD approximation of $f(x)=J_{1/2}(x+1)+1/(x^{2}+1)$ with truncation threshold $\epsilon$. The black squares correspond to $A_{n,m}=0$. The white lines mark when the total number of samples equals $n=80$ or $n^{\epsilon}\approx 54.8$, the numerical dimension. We conclude that $A_{n,m}^{\epsilon}$ accurately predicts the behaviour of the approximation error, while $A_{n,m}$ overestimates the amount of samples that are needed for accurate numerical approximation.
@@ 6.2 Approximating on irregular domains
Approximating a function on an irregular domain $D\subset\mathbb{R}^{d}$ is required for many computational problems, yet an orthonormal or Riesz basis is generally unknown for such a domain. Therefore, numerical methods often resort to working with an orthonormal basis defined on a bounding box surrounding $D$. Crucially, this set of functions does not form an orthonormal basis for the underlying domain $D$, but it does constitute an overcomplete frame, referred to as an extension frame. As discussed in section 2.1, a subsequence of an overcomplete frame inevitably becomes numerically redundant as its size increases and, hence, practical approximations are subject to $\ell^{2}$-regularization. Approximating on an irregular domain is, therefore, a typical scenario in which numerical redundancy arises.
@@
for odd $n$. Smooth functions, not necessarily periodic on $[-W,W]$, can be approximated in $V_{n}=\operatorname{span}(\Phi_{n})$ with exponential convergence [36]. The numerical convergence behaviour, taking into account the influence of regularization, is extensively analysed in [9].
In this section, we analyse how regularization affects the required amount of data for accurate approximation using the results from section 5, which focus on random pointwise sampling. The associated analysis outlined in section 5.3 depends heavily on a good understanding of the singular values and the singular vectors associated to the synthesis operator $\mathcal{T}_{n}$. For many spanning sets, little is known about this; the singular vectors of the Fourier extension frame are, however, related to discrete prolate spheroidal wave functions (DPSWFs), which have been analysed extensively in the field of signal processing. For a detailed explanation of this connection, see [40, section 3.1]. DPSWFs are an indispensable tool for bandlimited extrapolation, introduced and thoroughly analysed by Slepian and his collaborators during their time at Bell Labs [50]. In this setting, a notion of numerical dimension is not new, see e.g. [39] and [23, section 2.3]
We note that many high-dimensional approximation algorithms assume that the target function is defined on a tensor-product domain. A common example is surrogate model construction in uncertainty quantification. In practice, however, variables are often correlated, or the forward model may be infeasible on parts of the domain. As a result, the approximation effectively becomes a Fourier or polynomial frame approximation, as explained in [7]. To fully understand such methods, including their information-based complexity, a detailed analysis of the role of regularization is essential. The analysis in this section offers an initial step towards a deeper understanding of these approaches.
6.2.1 Numerical dimension
In the case of pointwise random sampling, the required number of samples for accurate regularized least squares approximation scales with $n^{\epsilon}\log(n^{\epsilon})$, as stated in Theorem 10. Here, $n^{\epsilon}$ denotes the numerical dimension. This theorem assumes that the samples are drawn randomly from a (near-)optimal sampling distribution defined by (50). The numerical dimension $n^{\epsilon}$ thus plays a similar role for regularized approximation as the dimension $n$ for unregularized approximation. The following theorem formalizes the dependence of $n^{\epsilon}$ on $n$, $W$ and $\epsilon$ for the one-dimensional Fourier extension frame.
@@ 6.2.1 Numerical dimension
such that the numerical dimension $n^{\epsilon}$ can be bounded by
@@
Note that $\sigma_{i}(\mathcal{T}_{n})^{2}$ equals the $i$-th eigenvalue of the Gram matrix $G_{n}$, representing the operator $\mathcal{T}_{n}^{*}\mathcal{T}_{n}$. Furthermore, $G_{n}$ is exactly the prolate matrix [38, eq. (1)]. Therefore, $\sigma_{i}(\mathcal{T}_{n})^{2}=\lambda_{i-1}$, for all $i=1,\dots,n$, where the latter is analysed in [38]. It follows from [38, Corollary 1] that (56) holds for (65). Moreover, (66) can be deduced by [38, Corollary 2]. The final result is a consequence of Theorem 11. ∎
A numerical verification of Theorem 12 is shown on Figure 8 for $n=141$ and $\epsilon=10\kern 1.0pt\operatorname{\epsilon_{\text{mach}}^{\text{sp}}}\approx 10^{-6}$ and $\epsilon=10\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}\approx 2\times 10^{-15}$, where $\operatorname{\epsilon_{\text{mach}}^{\text{sp}}}$ and $\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}$ denote the working precision associated with IEEE single and double precison floating point numbers, respectively. Note that the computation of $n^{\epsilon}$ is done in extended precision, such that the singular values $\sigma_{i}$ can be computed with a working precision significantly smaller than $\epsilon$. For implementational details, we refer to [32]. Both the theory and the numerical verification show that the numerical dimension essentially scales as $2nW$. This formalizes the intuition that increased redundancy, i.e., smaller $W$, results in a reduced amount of required information. Observe that $n^{\epsilon}=\mathcal{O}(n)$, such that asymptotically the number of samples needed for regularized and unregularized approximation is similar, namely $\mathcal{O}(n\log(n))$.
@@
Figure 8: The numerical dimension $n^{\epsilon}$ for the one-dimensional Fourier extension frame (64) as a function of the domain parameter $W$, for $n=141$. The dashed lines illustrate $n^{\epsilon}\leq n=141$ and the theoretical bound $n^{\epsilon}\leq\lceil 2nW\rceil+2+\frac{2}{\pi^{2}}\left(\log(8/\epsilon^{2})+1\right)\log(4n)$ stated in Theorem 12.
@@ 6.2.2 Uniformly random sampling
For many applications, one cannot freely choose the sample distribution. Another interesting question to answer in this regard is how many uniformly random samples are needed for accurate approximation. As follows from Theorems 9 and 10 for $w=1$, this number is proportional to the maximum of the inverse Christoffel function $k_{n}$ and the inverse numerical Christoffel function $k_{n}^{\epsilon}$ for unregularized and regularized least squares fitting, respectively. The following theorem shows that $\|k_{n}\|_{L^{\infty}[-W,W]}=\mathcal{O}(n^{2})$.
@@
For the spanning set defined by (64) with $W<1/2$, the Christoffel function satisfies
$\|k_{n}\|_{L^{\infty}[-W,W]}\geq k_{n}(W)\sim n^{2}\frac{\pi}{2}\cot(W\pi)\qquad\text{as $n\to\infty$}.$
@@
In [54] a related kernel is being analysed as a means to determine the pointwise convergence behaviour of Fourier extensions. More specifically, in [54], the kernel $K_{N}(x,y)$ is analysed, which we relate to the inverse Christoffel function $k_{n}$ in our setting via
$k_{n}(x)=K_{n}(x/W,x/W)/W,\qquad x\in[-W,W].$
We can, therefore, restrict ourselves to the analysis of $K_{N}(x,x)$ for $x\in[-1,1]$. The result can now be obtained by combining [54, Theorem 4.1] with [54, eq. 8], which is valid for large degree $n$ and for $x\in[1-\delta,1]$ for some $\delta>0$. After expanding all quantities for large $n$ and plugging in $x=1$, one obtains the asymptotic result via lengthy but straightforward calculations. The inequality $\|k_{n}\|_{L^{\infty}[-W,W]}\geq k_{n}(W)$ holds since $\|k_{n}\|_{L^{\infty}[-W,W]}=\|k_{n}\|_{\infty}$. ∎
The inverse numerical Christoffel function $k_{n}^{\epsilon}$ can be bounded using (59), which depends on the size of the plunge region $i_{2}-i_{1}$ and $\operatorname{S_{\text{tail}}}$. The latter is shown to be $\mathcal{O}(\log(n))$ in Theorem 12. The following theorem proves that the size of the plunge region is $\mathcal{O}(\log(n))$ as well.
@@
$\|k_{n}^{\epsilon}\|_{\infty}\leq\frac{n}{1-\epsilon^{2}}+\left(4+\frac{2}{\pi^{2}}\left(2\log\left(\frac{8}{\epsilon^{2}}\right)+1\right)\log(4n)\right)\max_{i_{1}<i\leq n}\|u_{i}\|_{\infty}^{2}.$ (69)
@@
Similarly as for the proof of Theorem 12, observe that $\sigma_{i}(\mathcal{T}_{n})^{2}=\lambda_{i-1}$, for all $i=1,\dots,n$, where the latter is analysed in [38]. It follows from [38, Corollary 1] that (55) holds for (68). By combining the results from Theorem 12 with (59) and (68), we arrive at (69). ∎
It remains to examine how the basis functions $u_{i}$ behave. As mentioned before, these functions are linked to the discrete prolate spheroidal wave functions. In Appendix D, we summarize known asymptotic results due to Slepian, which indicate that $\|u_{i}\|_{\infty}=\mathcal{O}(\sqrt{n})$ as $n\to\infty$ for all $i$, such that
$\|k_{n}^{\epsilon}\|_{\infty}=\mathcal{O}(n\log(n)).$
This behaviour is verified numerically in Figure 9, which confirms that $\|k_{n}\|_{\infty}=\mathcal{O}(n^{2})$, while $\|k_{n}^{\epsilon}\|_{\infty}$ appears to grow linearly in $n$. This difference translates in the need for log-linear or quadratic oversampling when computing a least squares approximation with or without regularization, respectively, using uniformly random samples.
@@
The approximation seems to converge root-exponentially. For a more detailed analysis of the convergence behaviour, we refer to [9]. Note that convergence is exponential without regularization, see e.g. [36] and the example in section 3, but demands quadratic oversampling. Hence, we get root-exponential convergence as a function of the number of sample points $m$ with and without regularization. This is an example of a setting where the negative effects of regularization, i.e., reduced accuracy, are balanced by the positive effects, i.e., reduced need for data.
@@
Remark 6.
What happens if we want to compute an approximation that converges root-exponentially all the way down to zero? In this case, we would need to vary the precision and the regularization parameter with $n$: $\epsilon\sim\epsilon_{\text{mach}}=\exp(-C\sqrt{n})$ for some $C>0$. Following Theorem 14, linear oversampling does not suffice in this case. Specifically, the required linear oversampling factor scales with $\log(1/\epsilon)$.
@@
Figure 9: The maximum of the inverse (numerical) Christoffel function for the one-dimensional Fourier extension frame (64) as a function of $n$. Full line: $\|k_{n}\|_{\infty}$, circles: $\|k_{n}^{\epsilon}\|_{\infty}$ for single precision $\epsilon=10\kern 1.0pt\operatorname{\epsilon_{\text{mach}}^{\text{sp}}}\approx 10^{-6}$, and squares: $\|k_{n}^{\epsilon}\|_{\infty}$ for double precision $\epsilon=10\kern 1.0pt\operatorname{\epsilon_{\text{mach}}^{\text{dp}}}\approx 2\times 10^{-15}$. The linear behaviour of $k_{n}^{\epsilon}$ versus the quadratic behaviour of $k_{n}$ implies a difference between log-linear and quadratic sampling rates for least squares fitting with uniformly random samples. The dashed lines mark $5\log_{10}(1/\epsilon)+C$, for varying constant $C$. They empirically match the behaviour of $k_{n}^{\epsilon}$.